OpenStudy (anonymous):

find the integral of sinxcosx using the identity sin(2x) = 2sinxcosx

OpenStudy (dumbcow):

\[\int\limits_{}^{}\sin x \cos x dx = \frac{1}{2}\int\limits_{}^{}\sin 2x dx\]

OpenStudy (anonymous):

how did you know to use sin? can you take me through step by step?

myininaya (myininaya):

also you don't have to use that identity let u=sinx du=cosx dx so we have \[\int\limits_{}^{}u du=\frac{u^2}{2}+C=\frac{(sinx)^2}{2}+C\]

myininaya (myininaya):

what do you mean how did you know to use sine? its says to in the problem and also sin(2x)=sin(x+x)=sinxcosx+sinxcosx=2sinxcosx

OpenStudy (anonymous):

oh, i didn't know how to evaluate it in the problem like that

OpenStudy (dumbcow):

yes that way is simpler, but if you have to use that identity then substitute sincos with 1/2 sin(2x) then use u substitution

myininaya (myininaya):

do you know substitution yet?

OpenStudy (anonymous):

yes....I completed the integration from your first answer dumbcow and i got the right answer. thanks everyone!

myininaya (myininaya):

have fun i love math too